OpenStudy (anonymous):

Identify the value of x that makes each pair of ratios equivalent 2:x and 12:18

OpenStudy (anonymous):

wait a minute >w>

OpenStudy (anonymous):

think i got it

OpenStudy (anonymous):

whats your answer?

OpenStudy (anonymous):

unsure about it but got 4

OpenStudy (anonymous):

wait no thats not right

OpenStudy (anonymous):

To find the answer to this ratio you will use cross-multiplication

OpenStudy (anonymous):

i know i got the right answer ^^

OpenStudy (anonymous):

3

OpenStudy (anonymous):

do you need anything else?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

ill help :)

OpenStudy (anonymous):

just needing 21 to x and 7 to 10 & x to 50 and 16 to 25 (the other test i got one wrong but did really well i was surprised with myself im grateful everyones helpful it made me abit better)

OpenStudy (anonymous):

Always happy to help :) Do you mean 21/x and 7/10 & x/50 and 16/25?

OpenStudy (anonymous):

it says 21 to x and 7 to 10 so im sure thats the same

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the other is listed the same way

OpenStudy (anonymous):

Cross-Multiply: 21/x and 7/10 21(10)=7x x/50 and 16/25 25x=50(16)

OpenStudy (anonymous):

i think somethings wrong ;o

OpenStudy (anonymous):

or im reading it wrong

OpenStudy (whpalmer4):

2:x = 12:18 for what value of x? 2/x = 12/18 if we cross multiply, 2*18 = 12*x 36 = 12*x x = 3 2:3 = 12:18 as a check, multiply the ratio on the left by 6 and verify that it equals the one on the right. 2*6:3*6 = 12:18 12:18 = 12:18 (multiplied by 6 because that's the ratio 12:2, just checking that the denominator of our answer also is in the ratio 12:2)

OpenStudy (whpalmer4):

21:x and 7:10 21*10 = 7*x 210 = 7x 30 = x x/50 as 16/25 x*25 = 50*16 x = 50*16/25 = 2*16 = 32

OpenStudy (anonymous):

;o whoaa

OpenStudy (anonymous):

ohh 30 i got 30

OpenStudy (whpalmer4):

do you still think something is wrong?